# The T2K Collaboration

### v1

@@ V.2.1 Reweighting of MC events
The weights were then organized into a set of splines, allowing for smooth interpolation between the node in a certain analysis bi, when the value of the dial was varied with respect to the nominal value. These splines allow for smooth interpolation between dial values, enabling continuous variation of model parameters. They constitute one of the inputs passed to the cross section fitter. This approach is the same as the one adopted in many other T2K cross section analyses  [11].
@@ V.2.2 CCQE and the axial mass
CCQE interactions are dominant at T2K energies. The CCQE cross section on a single nucleon depends on vector and axial form factors [7]. Vector form factors are known precisely from electron scattering data [32], due to the conserved vector current hypothesis. However, for the axial form factor the situation is different and it was a dominant source of uncertainty. A dipole expression is assumed with an axial mass parameter $M_{A}^{QE}=1.21$ GeV, which is constrained by past neutrino-deuterium scattering measurements [33].
In CCQE interactions on a nucleon, within a nucleus such as Oxygen or Carbon, the effects of nuclear binding were modeled using the Benhar Spectral Function (SF) model [34], which provides a realistic shell-model-based description of the nuclear ground state and is strongly favored over relativistic Fermi gas models [35] by electron scattering data [36].
@@ V.2.3 Multi-Nucleon Interactions (2p2h)
The term ‘2p2h process’ refers to neutrino interactions on pairs of correlated nucleons. They give rise to an important fraction of the events observed in T2K’s energy range. In NEUT simulations 2p2h includes contributions from Meson Exchange Currents, nucleon-nucleon correlations (NN) and their interference [37, 38] as described by the Valencia model [39]. It was assumed that the 2p2h contribution populates a kinematic region of momentum transfer $|\vec{q}|\leq 1.2$ GeV/c. Since precise 2p2h cross section measurements are unavailable, a conservative approach was taken in setting priors for its dials. A set of normalization dials scaling the whole 2p2h contribution (depending on the neutrino flavour and target nucleus) were considered in this analysis.
@@ V.2.4 Single Pion Production
Single-pion final states arise from resonance excitation, coherent pion production and multi-pion processes. As already mentioned in Sec. I, the neutrino flux observed at the off-axis angle of the WAGASCI-BabyMIND is shifted towards higher energies with respect to ND280 resulting in an increased probability of pion production from the neutrino interactions, thus making their correct characterization important for our analysis. The resonance excitation mechanism is modeled by NEUT with the Rein-Sehgal model [40, 41].
The most important dials were $M_{A}^{RES}$ (RESonance axial-vector mass), $C_{5}^{A}(0)$ (normalization of the leading axial form factor) and $I_{\nicefrac{{1}}{{2}}}$, (the amount of non-resonance contribution with isospin $\nicefrac{{1}}{{2}}$ in the Rein-Sehgal model). The nominal values were taken from theoretical models tuned to external data [42, 43].
@@ V.3 Detector systematics
Figure 13: Fractional errors due to the detector systematic uncertainties on the number of selected events. Subfigures (a) and (b) correspond to the WAGASCIs sample, while (c) and (d) correspond to the PM sample. The muon momentum binning is the same as in Tab. 6 and the muon angle binning is the same as in Tab. 7.
@@ VI Simulated data studies
The pseudo data were made with an alternate Random Phase Approximation model (BeRPA) for CCQE interactions instead of the default RPA Nieves model [44, 39].
@@
This alternative model accounts for the difference between data and MC reported by the MINERvA experiment [45] and the MINOS experiment [46]. The difference emerges in the low-energy transfer region (sub-GeV), where the data were smaller than the MC prediction.
@@ VII.3 Comparison with different interaction models
We used NEUT with a different choice of interaction models. One difference was in the treatments of the nuclear ground state. The alternative model described the ground state with a Global Fermi Gas (GFG) model [48]. The CCQE model was the Nieves 1p1h model without the RPA correction, with $M_{A}^{\mathrm{QE}}$ set to 1.03 GeV, compared to 1.21 GeV in the original NEUT model. The expected number of CCQE events decreased by about 15% in this model.
@@
GENIE allows for various combinations of interaction models to be simulated and tested. In this study the CCQE model was the same as NEUT but with a different $M_{A}^{\mathrm{QE}}$ (0.99 GeV). Final State Interactions were modeled with the hN model [49] and the treatment of the ground state of a nucleus was also different. The NEUT model utilized the Benhar SF model, whereas GENIE adopts a Relativistic Fermi Gas (RFG) model.
@@ Acknowledgements.
The T2K collaboration would like to thank the J-PARC staff for superb accelerator performance. We thank the CERN NA61/SHINE Collaboration for providing valuable particle production data. We acknowledge the support of MEXT, JSPS KAKENHI (JP16H06288, JP18K03682, JP18H03701, JP18H05537, JP19J01119, JP19J22440, JP19J22258, JP20H00162, JP20H00149, JP20J20304, JP24K17065) and bilateral programs (JPJSBP120204806, JPJSBP120209601), Japan; NSERC, the NRC, and CFI, Canada; the CEA and CNRS/IN2P3, France; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) 397763730, 517206441, Germany; the NKFIH (NKFIH 137812 and TKP2021-NKTA-64), Hungary; the INFN, Italy; the Ministry of Science and Higher Education (2023/WK/04) and the National Science Centre (UMO-2018/30/E/ST2/00441, UMO-2022/46/E/ST2/00336 and UMO-2021/43/D/ST2/01504), Poland; the RSF (RSF 24-12-00271) and the Ministry of Science and Higher Education, Russia; MICINN (PID2022-136297NB-I00 /AEI/10.13039/501100011033/ FEDER, UE, PID2021-124050NB-C31, PID2020-114687GB-I00, PID2019-104676GB-C33), Government of Andalucia (FQM160, SOMM17/6105/UGR) and the University of Tokyo ICRR’s Inter-University Research Program FY2023 Ref. J1, and ERDF and European Union NextGenerationEU funds (PRTR-C17.I1) and CERCA program, and University of Seville grant Ref. VIIPPIT-2023-V.4, and Secretariat for Universities and Research of the Ministry of Business and Knowledge of the Government of Catalonia and the European Social Fund (2022FI_B 00336), Spain; the SNSF and SERI (200021_185012, 200020_188533, 20FL21_186178I), Switzerland; the STFC and UKRI, UK; the DOE, USA; and NAFOSTED (103.99-2023.144,IZVSZ2.203433), Vietnam. We also thank CERN for the UA1/NOMAD magnet, DESY for the HERA-B magnet mover system, the BC DRI Group, Prairie DRI Group, ACENET, SciNet, and CalculQuebec consortia in the Digital Research Alliance of Canada, and GridPP in the United Kingdom, and the CNRS/IN2P3 Computing Center in France and NERSC (HEP-ERCAP0028625). In addition, the participation of individual researchers and institutions has been further supported by funds from the ERC (FP7), “la Caixa” Foundation (ID 100010434, fellowship code LCF/BQ/IN17/11620050), the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie grant agreement numbers 713673 and 754496, and H2020 grant numbers RISE-GA822070-JENNIFER2 2020 and RISE-GA872549-SK2HK; the JSPS, Japan; the Royal Society, UK; French ANR grant number ANR-19-CE31-0001 and ANR-21-CE31-0008; and Sorbonne Université Emergences programmes; the SNF Eccellenza grant number PCEFP2_203261; the VAST-JSPS (No. QTJP01.02/20-22); and the DOE Early Career programme, USA. For the purposes of open access, the authors have applied a Creative Commons Attribution license to any Author Accepted Manuscript version arising.

### v2

@@ V.2.1 Reweighting of MC events
The weights were then organized into a set of splines, allowing for smooth interpolation between the node in a certain analysis bin, when the value of the dial was varied with respect to the nominal value. These splines allow for smooth interpolation between dial values, enabling continuous variation of model parameters. They constitute one of the inputs passed to the cross section fitter. This approach is the same as the one adopted in many other T2K cross section analyses  [10.1093/ptep/ptab014].
@@ V.2.2 CCQE and the axial mass
CCQE interactions are dominant at T2K energies. The CCQE cross section on a single nucleon depends on vector and axial form factors [Llewellyn-Smith]. Vector form factors are known precisely from electron scattering data [PhysRevD.102.074012], due to the conserved vector current hypothesis. However, for the axial form factor the situation is different and it was a dominant source of uncertainty. A dipole expression is assumed with an axial mass parameter $M_{A}^{QE}=1.21$ GeV, which is constrained by past neutrino-deuterium scattering measurements [PhysRevD.104.093001].
In CCQE interactions on a nucleon, within a nucleus such as Oxygen or Carbon, the effects of nuclear binding were modeled using the Benhar Spectral Function (SF) model [Benhar:1999bg], which provides a realistic shell-model-based description of the nuclear ground state and is strongly favored over relativistic Fermi gas models [1975547] by electron scattering data [Sobczyk:2017mts].
@@ V.2.3 Multi-Nucleon Interactions (2p2h)
The term ‘2p2h process’ refers to neutrino interactions on pairs of correlated nucleons. They give rise to an important fraction of the events observed in T2K’s energy range. In NEUT simulations 2p2h includes contributions from Meson Exchange Currents, nucleon-nucleon correlations (NN) and their interference [RuizSimo:2016rtu, PASCHALIS2020135110] as described by the Valencia model [Nieves:2011pp]. It was assumed that the 2p2h contribution populates a kinematic region of momentum transfer $|\vec{q}|\leq 1.2$ GeV/c. Since precise 2p2h cross section measurements are unavailable, a conservative approach was taken in setting priors for its dials. A set of normalization dials scaling the whole 2p2h contribution (depending on the neutrino flavour and target nucleus) were considered in this analysis.
@@ V.2.4 Single Pion Production
Single-pion final states arise from resonance excitation, coherent pion production and multi-pion processes. As already mentioned in Sec. I, the neutrino flux observed at the off-axis angle of the WAGASCI-BabyMIND is shifted towards higher energies with respect to ND280 resulting in an increased probability of pion production from the neutrino interactions, thus making their correct characterization important for our analysis. The resonance excitation mechanism is modeled by NEUT with the Rein-Sehgal model [REIN198179, Rein:1987cb].
The most important dials were $M_{A}^{RES}$ (RESonance axial-vector mass), $C_{5}^{A}(0)$ (normalization of the leading axial form factor) and $I_{\nicefrac{{1}}{{2}}}$, (the amount of non-resonance contribution with isospin $\nicefrac{{1}}{{2}}$ in the Rein-Sehgal model). The nominal values were taken from theoretical models tuned to external data [Graczyk:2009qm, Kuzmin:2006dh].
@@ V.3 Detector systematics
Figure 13: Fractional errors due to the detector systematic uncertainties on the number of selected events. Subfigures (a) and (b) correspond to the WAGASCIs sample, while (c) and (d) correspond to the PM sample. The muon momentum binning is the same as in Tab. 6 and the muon angle binning is the same as in Tab. 7. Systematic uncertainties arising from the neutrino interaction model are not included in this plot, only detector and reconstruction-related systematic uncertainties are included.
@@ VI Simulated data studies
The pseudo data were made with an alternate Random Phase Approximation model (BeRPA) for CCQE interactions instead of the default RPA Nieves model [RPA-Default, Nieves:2011pp].
@@
This alternative model accounts for the difference between data and MC reported by the MINERvA experiment [MINERVA_pion_production] and the MINOS experiment [MINOS_ND_CCQE]. The difference emerges in the low-energy transfer region (sub-GeV), where the data were smaller than the MC prediction.
@@ VII.3 Comparison with different interaction models
We used NEUT with a different choice of interaction models. One difference was in the treatments of the nuclear ground state. The alternative model described the ground state with a Global Fermi Gas (GFG) model [NEUT_update]. The CCQE model was the Nieves 1p1h model without the RPA correction, with $M_{A}^{\mathrm{QE}}$ set to 1.03 GeV, compared to 1.21 GeV in the original NEUT model. The expected number of CCQE events decreased by about 15% in this model.
@@
GENIE allows for various combinations of interaction models to be simulated and tested. In this study the CCQE model was the same as NEUT but with a different $M_{A}^{\mathrm{QE}}$ (0.99 GeV). Final State Interactions were modeled with the hN model [GENIE_FSI] and the treatment of the ground state of a nucleus was also different. The NEUT model utilized the Benhar SF model, whereas GENIE adopts a Relativistic Fermi Gas (RFG) model.
@@ Acknowledgements.
The T2K collaboration would like to thank the J-PARC staff for superb accelerator performance. We thank the CERN NA61/SHINE Collaboration for providing valuable particle production data. We acknowledge the support of MEXT, JSPS KAKENHI (JP16H06288, JP18K03682, JP18H03701, JP18H05537, JP19J01119, JP19J22440, JP19J22258, JP20H00162, JP20H00149, JP20J20304, JP24K17065) and bilateral programs (JPJSBP120204806, JPJSBP120209601), Japan; NSERC, the NRC, and CFI, Canada; the CEA and CNRS/IN2P3, France; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) 397763730, 517206441, Germany; the NKFIH (NKFIH 137812 and TKP2021-NKTA-64), Hungary; the INFN, Italy; the Ministry of Science and Higher Education (2023/WK/04) and the National Science Centre (UMO-2018/30/E/ST2/00441, UMO-2022/46/E/ST2/00336 and UMO-2021/43/D/ST2/01504), Poland; the RSF (RSF 24-12-00271) and the Ministry of Science and Higher Education, Russia; MICINN (PID2022-136297NB-I00 /AEI/10.13039/501100011033/ FEDER, UE, PID2024-157541NB-I00 (UAM) and PID2023-146401NB-I00 (US), Severo Ochoa Centres of Excellence Programme 2025-2029 (CEX2024001442-S), Government of Andalucia (FQM160) and the University of Tokyo ICRR’s Inter-University Research Program FY2025 Ref. J1, and ERDF and European Union (UAM: H2020-MSCA-RISE-GA872549- SK2HK) and NextGenerationEU funds (PRTR-C17.I1) and Generalitat de Catalunya (AGAUR 2021-SGR-01506, CERCA program) University of Seville grant (RYC2022-035203-I funded by MICIU/AEI/10.13039/501100011033, ”ERDF a way of making Europe” and FSE+, Ayudas ”Atracción de Investigadores con Alto Potencial”. Ref. VIIPPIT-2025, and Secretariat for Universities and Research of the Ministry of Business and Knowledge of the Government of Catalonia and the European Social Fund (2022FI_B 00336), Spain; the SNSF and SERI (200021_185012, 200020_188533, 20FL21_186178I), Switzerland; the STFC and UKRI, UK; the DOE, USA; and NAFOSTED (103.99-2023.144,IZVSZ2.203433), Vietnam. We also thank CERN for the UA1/NOMAD magnet, DESY for the HERA-B magnet mover system, the BC DRI Group, Prairie DRI Group, ACENET, SciNet, and CalculQuebec consortia in the Digital Research Alliance of Canada, and GridPP in the United Kingdom, and the CNRS/IN2P3 Computing Center in France and NERSC (HEP-ERCAP0028625). In addition, the participation of individual researchers and institutions has been further supported by funds from the ERC (FP7), “la Caixa” Foundation (ID 100010434, fellowship code LCF/BQ/IN17/11620050), the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie grant agreement numbers 713673 and 754496, and H2020 grant numbers RISE-GA822070-JENNIFER2 2020 and RISE-GA872549-SK2HK, the Horizon Europe Marie Sklodowska-Curie Staff Exchange project JENNIFER3 grant 101183137; the JSPS, Japan; the Royal Society, UK; French ANR grant number ANR-19-CE31-0001 and ANR-21-CE31-0008; and Sorbonne Université Emergences programmes; the SNF Eccellenza grant number PCEFP2_203261; the VAST-JSPS (No. QTJP01.02/20-22); and the DOE Early Career programme, USA. For the purposes of open access, the authors have applied a Creative Commons Attribution license to any Author Accepted Manuscript version arising. Definitions ANR: French National Research Agency CEA: Commissariat à l’Energie Atomique (France) http://www.cea.fr CERCA: Centres de Recerca de Catalunya (Spain) https://cerca.cat/en/ CERN: European Organization for Nuclear Research (derived from the name ”Conseil Européen pour la Recherche Nucléaire”) CFI: Canada Foundation for Innovation (Canada) CNRS/IN2P3: Centre National de la Recherche Scientifique—Institut National de Physique Nucléaire et de Physique des Particules (France) http://www.in2p3.fr DFG: Deutsche Forschungsgemeinschaft [German Research Foundation] (Germany) http://www.dfg.de/en/ DESY: Deutsches Elektronen-Synchrotron, (Germany) DOE: Department of Energy (United States) https://energy.gov ERC: European Research Council (European Union) https://erc.europa.eu ERDF: European Regional Development Fund (European Union) http://ec.europa.eu/regional_policy/en/funding/erdf/ GridPP: Grid for Particle Physics https://www.gridpp.ac.uk HERA-B: Hadron-Elektron-Ring-Anlage-B, ”Hadron Electron Ring Facility for Bs“, http://www-hera-b.desy.de INFN: Istituto Nazionale di Fisica Nucleare [National Institute for Nuclear Physics] (Italy) http://home.infn.it/en/ J-PARC: Japan Proton Accelerator Research Complex (Japan) JSPS: Japan Society for the Promotion of Science (Japan) http://www.jsps.go.jp/english/ MES: Ministry of Education and Science http://dx.doi.org/10.13039/501100003443 (Russia) MEXT: Ministry of Education, Culture, Sports, Science and Technology (Japan) MICINN: Ministerio de Ciencia, Innovacion y Universidades, (Spain) http://www.ciencia.gob.es NA61: North Area experiment 61 (CERN) NERSC: National Energy Research Scientific Computing Center https://www.nersc.gov/ NCN: Narodowe Centre Nauki [National Science Centre] (Poland) https://www.ncn.gov.pl NII: National Institute of Informatics (Japan) http://www.nii.ac.jp/en/ NRC: National Research Council (Canada) NSERC: Natural Sciences and Engineering Research Council (Canada) NKFIH: National Research, Development and Innovation Office (Hungary) https://nkfih.gov.hu/palyazoknak RFBR: Russian Foundation for Basic Research, (Russia) Royal Society: http://dx.doi.org/10.13039/501100000288 (United Kingdom) https://royalsociety.org RSF: Russian Science Foundation (Russia) SciNet: (Canada) https://www.scinethpc.ca SINET4: Science Information Network 4 (Japan) http://w4a.sinet.ad.jp SNSF: Swiss National Science Foundation http://dx.doi.org/10.13039/501100001711 (Switzerland) http://www.snf.ch/en/Pages/default.aspx SERI: State Secretariat for Education, Research and Innovation http://dx.doi.org/10.13039/501100007352 (Switzerland) https://www.sbfi.admin.ch/sbfi/en/home.html STFC: Science and Technology Facilities Council, (United Kingdom) UKRI: UK Research and Innovation, (United Kingdom) https://www.ukri.org/ UA1: Underground Area experiment 1 Digital Research Alliance of Canada : https://alliancecan.ca ADDITIONAL INFO Alfred P. Sloan Foundation, FundRef ID http://dx.doi.org/10.13039/100000879 (United States/New York) Ministry of Education, Culture, Sports, Science and Technology, FundRef ID http://dx.doi.org/10.13039/501100001700 (Japan) Natural Sciences and Engineering Research Council of Canada, FundRef ID http://dx.doi.org/10.13039/501100000038 (Canada) National Research Council Canada, FundRef ID http://dx.doi.org/10.13039/501100000046 (Canada/Ontario) Canada Foundation for Innovation, FundRef ID http://dx.doi.org/10.13039/501100000196 (Canada/Ontario) Centre National de la Recherche Scientifique, FundRef ID http://dx.doi.org/10.13039/501100004794 (France) Commissariat à l’Énergie Atomique et aux Énergies Alternatives, FundRef ID http://dx.doi.org/10.13039/501100006489 (Republic of France) Deutsche Forschungsgemeinschaft, FundRef ID http://dx.doi.org/10.13039/501100001659 (Federal Republic of Germany) Department of Energy, FundRef ID http://dx.doi.org/10.13039/100000015 (United States/District of Columbia) European Regional Development Fund, FundRef ID http://dx.doi.org/10.13039/501100008530 (Kingdom of Belgium) Istituto Nazionale di Fisica Nucleare, FundRef ID http://dx.doi.org/10.13039/501100004007 (Repubblica Italiana) H2020 European Research Council, FundRef ID http://dx.doi.org/10.13039/100010663 (Kingdom of Belgium) Narodowe Centrum Nauki, FundRef ID http://dx.doi.org/10.13039/501100004281 (Republic of Poland) Ministerstwo Nauki i Szkolnictwa Wy?szego, FundRef ID http://dx.doi.org/10.13039/501100004569 (Republic of Poland) Ministry of Education and Science, FundRef ID http://dx.doi.org/10.13039/501100003443 (Russia) Ministerio de Economía y Competitividad, FundRef ID http://dx.doi.org/10.13039/501100003329 (Kingdom of Spain) Russian Science Foundation, FundRef ID http://dx.doi.org/10.13039/501100006769 (Russian Federation) Russian Foundation for Basic Research, FundRef ID http://dx.doi.org/10.13039/501100002261 (Russian Federation) Royal Society, FundRef ID http://dx.doi.org/10.13039/501100000288 (United Kingdom of Great Britain and Northern Ireland) Science and Technology Facilities Council, FundRef ID http://dx.doi.org/10.13039/501100000271 (United Kingdom of Great Britain and Northern Ireland) State Secretariat for Education, Research and Innovation, FundRef ID http://dx.doi.org/10.13039/501100007352 (Switzerland) Swiss National Science Foundation, FundRef ID http://dx.doi.org/10.13039/501100001711 (Switzerland) Canada Foundation for Innovation (Canada) https://www.innovation.ca Japan Society for the Promotion of Science (Japan) http://www.jsps.go.jp/english/